# SMARTER: A Data‑efficient Framework to Improve Toxicity Detection with Explanation via Self-augmenting Large Language Models

### v2

@@ 9 Limitations
Due to budget constraints, we are only able to human validation for 1 dataset. In practice, it is recommended to solicit a larger pool of annotators to ensure appropriate amount of human oversight.
@@ 10 Ethical Considerations
Appendix A Amazon Turks Mechanical Annotation
We obtained IRB approval before soliciting annotation from the crowd workers. 14 workers were recruited to annotate the 342 explanations introduced in subsection 6.1. Samples are assigned at random by the AMT platform to each annotator. Due to budget constraint, we opt to collect annotation for HateXplain to preserve statistical power instead of spreading across all 3 datasets. The breakdown of the annotators’ demographic is below:
@@ 64 DPO 3 1e-04 3 1e-06
Implicit Hate $\diamond$ White Grievance: includes frustration over a minority groups perceived privilege and casting majority groups as the real victims of racism. This language is linked to extremist behavior and support for violence. $\diamond$ Inferiority Language: implies one group or individual is inferior to another, and it can include dehumanization, denial of a person’s humanity, and toxic language that compares the target with disease, insects, animals. Related to assaults on human dignity, dominance, and declarations of superiority of the in group. $\diamond$ Irony: refers to the use of sarcasm , humor, and satire to attack or demean a protected class or individual. $\diamond$ Stereotypes And Misinformation: associate a protected class with negative attributes such as crime, or terrorism. includes misinformation that feeds stereotypes and vice versa, like holocaust denial and other forms of historical negationism. $\diamond$ Threatening And Intimidation: conveys a speaker’s commitment to a target’s pain, injury, damage, loss or violation of rights, threats related to implicit violation of rights and freedoms, removal of opportunities, and more subtle forms of intimidation.
@@ Appendix G Examples of Explanations
We provide some examples of the explanations generated during our cross-training phrase in Figure 12, Figure 13, and Figure 14.

### v3

@@ 9 Limitations
Augmentation to base framework
In our experiments, samples are randomly sampled from the label space. However, more strategic sampling choices (e.g., selecting based on how the level of classification difficulty calibrated on validation set) may boost performance. We encourage practitioners to explore augmentation to maximize the utility of our framework.
@@
Due to budget constraints, we are only able to perform human validation for 1 dataset. In practice, it is recommended to solicit a larger pool of annotators to ensure appropriate amount of human oversight.
@@ 10 Ethical Considerations
11 Acknowledgment
This work is funded by the NSF under Grant No. 2229885 (NSF Institute for Trustworthy AI in Law and Society, TRAILS). We thank the service of ACL ARR reviewers, area chairs and the editors of the ACL conference for our paper’s publication.
@@ 11 Acknowledgment
Appendix A Amazon Mechanical Turk Annotation
We obtained IRB approval before soliciting annotation from the crowd workers. 14 workers were recruited to annotate the 342 explanations introduced in Section 6.1. Samples are assigned at random by the AMT platform to each annotator. Due to budget constraint, we opt to collect annotation for HateXplain to preserve statistical power instead of spreading across all 3 datasets. The breakdown of the annotators’ demographic is below:
@@ Appendix C Obtaining Explanations for In-domain Data
Dataset Model Min Mean Max Range
HateXplain GPT-4o-mini 0.22 0.29 0.42 0.20
GPT-4.1 0.47 0.52 0.60 0.13
GPT-5-chat 0.61 0.62 0.63 0.02
Qwen-32B 0.52 0.55 0.59 0.07
Latent Hate GPT-4o-mini 0.17 0.25 0.34 0.17
GPT-4.1 0.58 0.63 0.67 0.09
GPT-5-chat 0.54 0.60 0.65 0.11
Qwen-32B 0.54 0.57 0.60 0.06
Implicit Hate GPT-4o-mini 0.14 0.15 0.16 0.02
GPT-4.1 0.30 0.38 0.51 0.21
GPT-5-chat 0.29 0.40 0.51 0.22
Qwen-32B 0.44 0.48 0.51 0.07
Table 6: Detailed 16-shot in-context learning (ICL) results for commercial models across 3 random seeds. For each dataset, we report the minimum, mean, and maximum Macro F1 scores, along with the range (Max – Min) to illustrate performance variability. Models were prompted with 16 examples per class. High variance, particularly for GPT-4o-mini and on complex tasks like Implicit Hate, demonstrates ICL’s brittleness for nuanced multi-class toxicity detection.
@@ 64 DPO 3 1e-04 3 1e-06
Implicit Hate $\diamond$ White Grievance: includes frustration over a minority groups’ perceived privilege and casting majority groups as the real victims of racism. This language is linked to extremist behavior and support for violence. $\diamond$ Inferiority Language: implies one group or individual is inferior to another, and it can include dehumanization, denial of a person’s humanity, and toxic language that compares the target with disease, insects, animals. Related to assaults on human dignity, dominance, and declarations of superiority of the in group. $\diamond$ Irony: refers to the use of sarcasm , humor, and satire to attack or demean a protected class or individual. $\diamond$ Stereotypes And Misinformation: associate a protected class with negative attributes such as crime, or terrorism. includes misinformation that feeds stereotypes and vice versa, like Holocaust denial and other forms of historical negationism. $\diamond$ Threatening And Intimidation: conveys a speaker’s commitment to a target’s pain, injury, damage, loss or violation of rights, threats related to implicit violation of rights and freedoms, removal of opportunities, and more subtle forms of intimidation. $\diamond$ Incitement to Violence: includes flaunting in group unity and power or elevating known hate groups and ideologies
@@ Appendix G Examples of Explanations
We provide some examples of the explanations generated during our cross-training phase in Figure 13, Figure 14, and Figure 15.